A small canteen offers special meal package to their customers. The profit P that the canteen will earn at x number of customer is modeled by the function P(x) = -25x2 + 100x - 3000. How many people are needed for a profit of at least
Php5,000

Found 2 solutions by ankor@dixie-net.com, ikleyn:
Answer by ankor@dixie-net.com(22740)

A small canteen offers special meal package to their customers.
The profit P that the canteen will earn at x number of customer is modeled by the function P(x) = -25x^2 + 100x - 3000.
How many people are needed for a profit of at least Php5,000
:
There is something wrong with this equation, no matter what value you have for x, it will yield a negative answer.
;
An equation of 25x^2 + 100x - 3000 will work, a profit will be shown when there are 10 people or more
Using this to find the number required for an 5000 profit
-25x^2 + 100x - 3000 = 5000
-25x^2 + 100x - 3000 - 5000 = 0
-25x^2 + 100x - 8000
simplify, divide by 25
x^2 + 4x - 320 = 0
factors to
(x+20)(x-16) = 0
the positive solution is all we want here
x = 16 people will give a profit of 5000

Answer by ikleyn(52872)

.
A small canteen offers special meal package to their customers. The profit P that the canteen
will earn at x number of customer is modeled by the function P(x) = -25x2 + 100x - 3000.
How many people are needed for a profit of at least Php5,000
~~~~~~~~~~~~~~


To answer the question, you need to solve this equation

    -25x^2 + 100x - 3000 = 5000.


Simplify it step by step

    -25x^2 + 100x - 8000 = 0

     25x^2 - 100x + 8000 = 0

       x^2 -   4x + 320 = 0


The discriminant is  b^2 - 4ac = 4^2 - 4*320, which is negative value.


Hence, the given equation HAS NO real roots.


ANSWER.   The problem, as it is worded, printed, posted and presented is a FAKE.

The problem is a  FAKE.


By posting a fake, you steal my time.


-------------

A special note to tutor @ankor@dixie-net.com

    In such circumstances, a tutor should not edit the post of the visitor 

    and should not edit his (or her) numerical input data / numbers.



    The necessary and sufficient action is to inform the visitor about the discrepancy and to issue a WARNING MESSAGE.


    Providing the correct input is ENTIRELY the job and the responsibility of the visitor.